OpenStudy (anonymous):

Write the slope-intercept form of the equation of the line parallel to y = 5x + 2 which passes through the point (4, 9). A. y = 5x - 11 B. y = -5x + 9 C. y = -(1/5)x - 11

OpenStudy (hoblos):

parallel lines have same slope so the slope is 5 y-9=5(x-4) y=5x-11